# Testing Bell inequalities and probing quantum entanglement at a muon collider

### v1

@@ I   Introduction
Quantum entanglement [1] is one of the most distinctive and counter-intuitive features of quantum mechanics. Particles that have interacted in the past remain in an entangled state even when they are spatially separated. Particles in an entangled state exhibit correlations in their physical properties, such as spin polarization. A correlated quantum system may even lead to the violation of the Bell inequalities [2], which implies that the correlations cannot be explained by local hidden variables. Quantum entanglement is most readily observable in two-qubit systems, such as a pair of spin-1/2 particles or a pair of photons [3, 4, 5, 6]. Experimental searches for quantum entanglement and violation of Bell inequalities have been successfully performed in two-outcome measurements with correlated photon pairs [7, 8]. Other proposals have been made to test Bell inequalities in $e^{+}e^{-}$ collision events [9], in charmonium decays [10, 11], and more recently in $t\bar{t}$ events at hadron colliders  [12, 13, 14, 15, 16, 17, 18].
@@
The study of the quantum entanglement and testing Bell inequality violation can be another new subject for the high-energy physics community. Detectors at the high energy colliders, while not designed for the probing of entanglement, turn out to be surprisingly good in performing this task, thus initiating possibilities of interesting new measurements for quantum information science as well as physics discoveries beyond the Standard Model. Recently, spin polarization of pairs of top and anti-top quarks has been measured and entanglement between them are observed in $t\bar{t}$ events at the Large Hadron Collider (LHC) [19, 2024hgo, 20] . It has also been shown that Bell inequality is violated in the decays of B mesons at LHCb and Belle II experiment [21]. Interestingly, there are many new proposals made to test violation of an optimized Bell inequality and CGLMP (Collins-Gisin-Linden-Massar-Popescu) inequality [22] for two-qutrit system, using gauge boson production from Higgs boson decay [23, 24, 25, 26, 27, 28, 29, 30], heavy lepton pair production at a lepton collider and other future collider [31, 32, 33], and also from vector boson scattering process [34].
@@
Massive gauge bosons may provide an efficient way to explore quantum entanglement and test the CGLMP inequality because they play as a polarimeter of their own (for W boson) and the higher dimensionality of their entangled system. The corresponding spin polarizations can of gauge boson can be reconstructed from the angular distribution of the final leptons or jets. It has been argued that the observation of quantum entanglement and testing Bell inequality could be achieved in di-boson production in proton-proton collision as well as a lepton collider [35] with relatively better significance through a quantum state tomography technique that provides the density matrix of the entangled two-qutrit system.
@@
In this paper, we examine probing quantum entanglement and violation of Bell inequality at a Muon Collider through a detailed MC simulation. The Muon Collider recently receives revived interest and promises to be able to extend the Center-Mass (CM) energy to 1-10 TeV range or above, thus producing processes hard to observe at the current CM energy of the LHC, such as double Higgs boson production through vector boson fusion, top quark pair production in association with scalar and gauge boson, diboson production etc [36, 37, 38, 39].
@@ II.1 Density matrix and observable quantity
Defining the z-axis as the direction of on-shell Z boson’s 3-momentum, we can choose $J_{z}$ as the polarization operator and eigenstates of $J_{z}$ as the basis of the spin space. In our case with two spin-1 massive bosons’ system (i.e., two-qutrit system), both Z bosons are from Higgs decay, and the spin component is conserved in the CM frame. Therefore, the ZZ state can only lie in one of 3 joint states, i.e., ${\left|l_{1}l_{2}\right\rangle\in\{\left|-+\right\rangle,\left|00\right\rangle,\left|+-\right\rangle}\}$, where $l_{1}$ and $l_{2}$ are the spin states of two Z bosons. In quantum mechanics, the density matrix is a useful tool to describe the state of a quantum system. By definition, the density matrix of a two-qutrit ZZ system can be written as a tensor production
@@
where $p_{l_{1}l_{2}}$ is the possibility on $\left|l_{1}l_{2}\right\rangle$, and $p_{l_{1}l_{2}}\geq 1,\quad\sum p_{l_{1}l_{2}}=1$ [25].
@@
For more general case, we can construct the wave vector $\varepsilon^{\mu}(p,\lambda)$, which is the polarization vector of a spin-1 particle with momentum $p$ and helicity $\lambda$. In our case, we choose Minkowski metric $g_{\mu\nu}=\mathrm{diag}(1,-1,-1,-1)$. For the benefit of the relative simplicity of our expression, we first boost the Z boson to the CM frame of the Higgs boson. Then we determine z-axis which z unitary vector $\hat{k}$ as the direction of the on-shell Z boson, and determine XOZ plane which contains $\hat{k}$ and unit vector $\hat{x}$ in laboratory frame, so that we can write x-axis unit vector $\hat{r}=\mathrm{sign}(\cos\Theta)(\hat{x}-\hat{k}\cos\Theta)/\sin\Theta$. At present, we have constructed the z-axis and x-axis unit vectors and easily get y-axis $\hat{n}=\hat{k}\times\hat{r}$ to obtain a 3 dim reference system frame as figure 1. On this basis, we will construct the boosted basis vector in the lepton CM frame according to the coordinate system of the 3-dim. The boosted basis vector can be expressed as
$n_{1}^{\mu}(1)=n_{1}^{\mu}(2)=(0,\hat{n}),~{}~{}~{}n_{2}^{\mu}(1)=n_{2}^{\mu}(2)=(0,\hat{r}),~{}~{}~{}n_{3}^{\mu}(1)=\gamma(\beta,\hat{k}),~{}~{}~{}n_{3}^{\mu}(2)=\gamma(-\beta,\hat{k}),$ (2)
where $-\beta,\gamma$ is CM frame Lorentz boost parameters, and the indices $1,2$ represent the two Z bosons. The polarization vector can be written by the representation of eigenstates of $S_{3}$ as
@@
The probability amplitude of the $ZZ$ process is then formulated as
@@
which is the probability amplitude for $Z_{1}$ with momentum $p_{1}$ and helicity $\lambda_{1}$, and $Z_{2}$ with momentum $p_{2}$ and helicity $\lambda_{2}$. By the definition of the density matrix, in general, the density matrix of the two-qutrit system can be written as $\rho=\sum p_{ll^{\prime}}\left|l\right\rangle\left\langle l^{\prime}\right|$ According to the discussion above, while $\left|l\right\rangle$ is the state function of the two $Z$ bosons, this joint state function is actually
$\left|\Phi\right\rangle=\sum c_{ij}\left|ij\right\rangle\longrightarrow\sum\mathcal{M}(\lambda_{1},\lambda_{2})\left|\lambda_{1},\lambda_{2}\right\rangle,$ (5)
where $c_{ij}$ is the coefficient of the joint state function. Therefore, the density matrix components is
$\rho=\left|\Phi\right\rangle\left\langle\Phi\right|\rightarrow\rho_{iji^{\prime}j^{\prime}}$ $=c_{ij}c_{i^{\prime}j^{\prime}}^{*}$ (6)
$=\frac{1}{|\bar{\mathcal{M}}|^{2}}\mathcal{M}(\lambda_{1},\lambda_{2})\mathcal{M}(\lambda_{1}^{\prime},\lambda_{2}^{\prime})^{*}.$ (7)
Now we can represent density matrix components by probability amplitude defined in Eq. 4 as
$\rho\left(\lambda_{1},\lambda_{1}^{\prime},\lambda_{2},\lambda_{2}^{\prime}\right)=\frac{\mathcal{M}(\lambda_{1},\lambda_{2})\mathcal{M}(\lambda_{1}^{\prime},\lambda_{2}^{\prime})}{|\bar{\mathcal{M}}|^{2}}=\frac{1}{|\bar{\mathcal{M}}|^{2}}\left(\mathcal{M}_{\mu\nu}\varepsilon^{\mu}\left(p_{1},\lambda_{1}\right)\varepsilon^{\nu}\left(p_{2},\lambda_{2}\right)\right)\left(\mathcal{M}_{\mu^{\prime}\nu^{\prime}}\varepsilon^{\mu^{\prime}}\left(p_{1},\lambda_{1}^{\prime}\right)\varepsilon^{\nu^{\prime}}\left(p_{2},\lambda_{2}^{\prime}\right)\right)^{{\dagger}}.$ (8)
Here $\bar{\mathcal{M}}$ is the normalization factor of the probability amplitude. Moreover, we can construct the covariance helicity projector operator $\mathscr{P}_{\lambda\lambda^{\prime}}^{\mu\nu}$from Eq. 3 in the lepton CM frame boosted by momentum $p$, which can be defined as
$\mathscr{P}_{\lambda\lambda^{\prime}}^{\mu\nu}\left(p\right)$ $=\varepsilon^{\mu}\left(p,\lambda\right)\varepsilon^{\nu}\left(p,\lambda^{\prime}\right).$ (9)
Using the covariance helicity projector operator, the density matrix can be simplified as
@@
The form of the density matrix describing the polarization state of the two-qutrit system formed by two spin-1 bosons can generally be parameterized using Gell-Mann matrices  [35, 42]
$\rho(\lambda_{1},\lambda_{1}^{\prime},\lambda_{2},\lambda_{2}^{\prime})=\Big{(}\frac{1}{9}\left[\mathbb{1}_{3}\otimes\mathbb{1}_{3}\right]+\sum_{a}f_{a}\left[T^{a}\otimes\mathbb{1}_{3}\right]+\sum_{a}g_{a}\left[\mathbb{1}_{3}\otimes T^{a}\right]+\sum_{ab}h_{ab}\left[T^{a}\otimes T^{b}\right]\Big{)}_{\lambda_{1}\lambda_{1}^{\prime}\lambda_{2}\lambda_{2}^{\prime}},$ (11)
where $T^{a}$, $T^{b}$ are the $3\times 3$ Gell-Mann matrices with $a=1,2,\dots,8$. The coefficients $f_{a}$ and $g_{a}$ are components of the spin polarization vector, and $h_{ab}$ being correlation coefficients for $ZZ$ system. In this way, we will have 80 parameters to be determined from observed data. Applications of this method to probe quantum entanglement can be found in the papers [35, 42] Using Gell-Mann matrices to represent the density matrix 11 is one of the possible parameterization. There is another simple yet effective way to parameterize the density matrix by using irreducible tensor operators [43, 44, 45]
@@
where $T^{L}_{M}$ are the irreducible tensor operators complying with a trace rule such that Tr[${\rm Tr}\left[T_{M}^{L}(T^{L}_{M})^{\dagger}\right]$ = 3. It should be noted that a sum over the indices $L=1,2$ and $-L\leq M\leq L$. These tensor operators are usually expressed as linear combination of the spin spin operators for spin-1 particles. One may refer to paper [25] for the exact form of these operators. The coefficients $A^{i}_{LM}$ and $C_{L_{1}M_{1}L_{2}M_{2}}$ are similar to the ones in Eq. 11 representing both the polarizations and spin correlations. In both parametrization form, we may extract the corresponding coefficients from simulated or actual data, which is very challenging because both the CM frame of the colliding beam and the rest frame of the decaying particles should be established correctly so that the angular distributions of the final lepton can be determined precisely. Consequently, there will be multiple Lorentz boost or transformation between coordinate systems. However, in the tensor-parametrization form of the density matrix, the number of parameters to be determined reduces to three parameters. In order to extract the coefficients of the density matrix, a relation between these coefficients and a direct observable in the experiment is needed. This has been done through a quantum state tomography for weak decays of massive bosons [46]. The high energy collider experiments directly provide the general observable quantity, cross section, which can be expressed in terms of decaying matrix density of the massive bosons. For example, the differential cross section fro $ZZ\to\ell_{1}^{+}\ell_{1}^{-}\ell_{2}^{+}\ell_{2}^{-}$ can be written [45]
@@
where the solid angles $d\Omega^{\pm}\sin\theta^{\pm}d\theta^{\pm}d\phi^{\pm}$ are given in terms of the spherical coordinates for the momenta of the final charged leptons with respect to the rest frame of the decaying Z boson. $\rho_{V_{1}V_{2}}$ is the density matrix for the entangled $ZZ$ state. The decaying density matrix of a Z boson into charged leptons is given as [45, 25, 24]
@@
Now the coefficients in Eq. 12 can be expressed using the above differential cross section making use of orthogonal property of the spherical harmonics:
@@
It is worth noting that the $ZZ$ system is in the singlet state because the third component of the spin along the boson momentum direction is conserved. This imposes strong constraints on the density matrix, including only nine non-zero elements with the relation
@@ II.2 Quantum entanglement and Bell inequality for two-qutrit system
To construct the optimized Bell inequality, two observers A and B, each having two measurements: $A_{1}$ and $A_{2}$ for $A$, and $B_{1}$ and $B_{2}$ for $B$, are considered. For each value of $A$ and $B$, there are three possible outcomes. One can denote by $P(A_{i}=B_{j}+k)$ the probability that the outcomes $A_{i}$ and $B_{j}$ differ by $k$ modulo 3. Then we can obtain a simple function by linear sum of these functions:
@@
Once we have these coefficients determined from data, one can immediately calculate the density matrix for the composite system, allowing to probe the entanglement and test Bell inequality violation through the expectation values of the Bell operator.
@@
where $\mathcal{B}$ is the quantum Bell operator [47]. To test the above inequality in quantum mechanics we need to calculate the expectation value of the Bell operator given in [25] as
@@
For the two-qutrit systems where entanglement occurs, we can compute the corresponding expectation values of these operators and compare the larges of them. It is worth noting that the maximal violation of  20 obtained with the expectation values of the above Bell operators might not occur because we choose the Z axis of the outgoing boson as the polarization axis. However, a unitary rotation that maximizes the entanglement in the spin space formed from the spin basis is useful so that one maybe able to observe maximal violation of  20. This requires a unitary transformation on the above Bell operators as
where U and V are the independent three-dimensional unitary matrices. This optimization or maximization is done bin-by-bin in the following Monte-Carlo analysis and we give the form of these unitary matrices parameterised the spherical coordinates in the gauge boson rest frame.

### v2

@@ I   Introduction
Quantum entanglement [1] is one of the most distinctive and counter-intuitive features of quantum mechanics. Particles that have interacted in the past remain in an entangled state even when spatially separated. Particles in an entangled state exhibit correlations in their physical properties, such as spin polarization. A correlated quantum system may even lead to the violation of the Bell inequalities [2], which implies that the correlations cannot be explained by local hidden variables. Quantum entanglement can be readily observed in two-qubit system, such as a pair of spin-1/2 particles or a pair of photons. Experimental searches for quantum entanglement and violation of Bell inequalities have been successfully performed in two-outcome measurements with correlated photon pairs [3, 4, 5]. Other proposals have been made to test Bell inequalities in $e^{+}e^{-}$ collision events [6], in charmonium decays [7, 8], and more recently in $t\bar{t}$ events at hadron colliders  [9, 10, 11, 12, 13, 14, 15, 16].
@@
The study of the quantum entanglement and testing Bell inequality violation can be another new subject for the high-energy physics community. While detectors at high-energy colliders are not specifically designed to probe quantum entanglement, they have demonstrated surprising effectiveness in this task. This opens up exciting opportunities for novel measurements in quantum information science, as well as potential discoveries that could extend beyond the Standard Model. Recently, spin polarization of pairs of top and anti-top quarks has been measured and quantum entanglement is observed [17, 18, 19]. It has also been shown that Bell inequality is violated in the decays of B mesons at LHCb and Belle II experiment [20]. Interestingly, several new proposals made to test violation of Bell inequality, such as CGLMP (Collins-Gisin-Linden-Massar-Popescu) inequality [21], an optimized one for two-qutrit system. These proposals leverage gauge boson production from Higgs boson decay [22, 23, 24, 25, 26, 27, 28, 29], heavy lepton pair production at a lepton collider and other future collider [30, 31, 32], and also from vector boson scattering process [33].
@@
Massive gauge bosons may provide an efficient way to explore quantum entanglement and test the CGLMP inequality because they can act as their own polarimeters (in case of W boson) and exhibit higher dimensionality in their entangled system. The corresponding spin polarizations can be reconstructed from the angular distribution of the final leptons or jets. It has been argued that the observation of quantum entanglement and testing Bell inequality could be achieved in di-boson production in proton-proton collision as well as a lepton collider with relatively better significance by determining the polarization density matrix of the entangled two-qutrit system [34].
@@
In this paper, we investigate quantum entanglement and violation of Bell inequality, the CGLMP inequality, at a Muon Collider through detailed Monte Carlo (MC) simulation. The Muon Collider recently received revived interest due to its potential to achieve Center of Mass (CM) energies to 1-10 TeV range or beyond. Thus a Muon Collider allows for the observation of processes that are challenging to detect at the current CM energy of the LHC. These processes may include double Higgs boson production through Vector Boson Fusion (VBF), top quark pair production in association with scalar and gauge boson, diboson production etc [35, 36, 37, 38].
@@ II.1 Density matrix and observable quantity
The density matrix for the joint $ZZ$ system, $\rho$ acts on the 9 dimensional Hilbert space defined by the three spin states of each Z boson. Defining the z-axis as the direction of Z boson’s 3-momentum, we can choose the eigenstates of the momentum operator $J_{z}$ as the basis vectors of the spin space. The $Z$ boson pairs are produced from Higgs decay, and the spin component is conserved in the momentum direction in the CM frame. Therefore, the ZZ state can only lie in one of 3 joint states, i.e., ${\left|l_{1}l_{2}\right\rangle\in\{\left|-+\right\rangle,\left|00\right\rangle,\left|+-\right\rangle}\}$, where $l_{1}$ and $l_{2}$ are the spin states of two Z bosons. By definition, the density matrix of a two-qutrit ZZ system can be written as a tensor production
@@
where $p_{l_{1}l_{2}}\geq 0$ and $\quad\sum p_{l_{1}l_{2}}=1$.
@@
Without the loss of generality, we choose to work in the CM frame, where the unit vectors and corresponding reference system is shown in Fig1. 1. The spin eigenstates of each Z boson is defined in their respective rest frame. Therefore, we need to boost $Z$ boson into the CM frame of the Higgs boson. The reference system in which the angular distribution of the decay leptons are determined is similar to the one applied in tp pair production [41] and massive gauge boson production in Higgs decay [34, 24]. Firs of all, a unit vector $\hat{z}$ is defined in the three-momentum direction of the on-shell $Z$ boson in the Higgs boson rest frame. Secondly, the $\hat{x}$ vector on the production plane is defined as $\hat{x}={\rm sign}(\cos\theta)(\hat{p}_{V}-\hat{z}\cos\theta)/\sin\theta$, with $\hat{p_{V}}=(0,0,1)$ the direction of the incident muon beams in the laboratory frame. Finally, the unit vector $\hat{y}$ is defined as $\hat{y}=\hat{z}\times\hat{x}$, perpendicular to the production plane.
The polarization vector of the spin-1 particle can be expressed as a linear combination of the three reference vectors as
@@
where the boosted vectors $n_{i}^{\mu}$ for each spin-1 particle are defined as
$n_{1}^{\mu}(1)=n_{1}^{\mu}(2)=(0,\hat{y}),~{}~{}~{}n_{2}^{\mu}(1)=n_{2}^{\mu}(2)=(0,\hat{x}),~{}~{}~{}n_{3}^{\mu}(1)=\gamma(\beta,\hat{z}),~{}~{}~{}n_{3}^{\mu}(2)=\gamma(-\beta,\hat{z}),$ (3)
with $-\beta,\gamma$ boost parameters, and the indices $1,2$ represent the two Z bosons. The probability amplitude of the $ZZ$ productin with helicities $\lambda_{1}$ and $\lambda_{2}$ is given by
@@
The, the polarization density matrix of the $ZZ$ joint system can be written using the helicity basis as
$\rho=\left|\Phi\right\rangle\left\langle\Phi\right|$ $=\frac{1}{|\bar{\mathcal{M}}|^{2}}\mathcal{M}(\lambda_{1},\lambda_{2})\mathcal{M}(\lambda_{1}^{\prime},\lambda_{2}^{\prime})^{*}$ (5)
$=\frac{1}{|\bar{\mathcal{M}}|^{2}}\left(\mathcal{M}_{\mu\nu}\varepsilon^{\mu}\left(p_{1},\lambda_{1}\right)\varepsilon^{\nu}\left(p_{2},\lambda_{2}\right)\right)\left(\mathcal{M}_{\mu^{\prime}\nu^{\prime}}\varepsilon^{\mu^{\prime}}\left(p_{1},\lambda_{1}^{\prime}\right)\varepsilon^{\nu^{\prime}}\left(p_{2},\lambda_{2}^{\prime}\right)\right)^{{\dagger}}.$ (6)
where $\bar{\mathcal{M}}$ is the spin-averaged unpolarized scattering amplitude. Using the polarization vector, we can construct the covariant helicity projection operator $\mathscr{P}_{\lambda\lambda^{\prime}}^{\mu\nu}$as
$\mathscr{P}_{\lambda\lambda^{\prime}}^{\mu\nu}\left(p\right)$ $=\varepsilon^{\mu}\left(p,\lambda\right)\varepsilon^{\nu}\left(p,\lambda^{\prime}\right),$ (7)
and finally the density matrix for the joint system reads as
@@
To obtain the density matrix from simulated or actual data, a proper form of parametrization is needed. The form of the density matrix describing the polarization state of the two-qutrit system formed by two spin-1 bosons can generally be parameterized using $3\times 3$ matrices composed of either Gell-Mann matrices  [34, 42] or spin-1 component operators. Using Gell-Mann matrices to represent the density matrix is one of the possible parameterization. There is another simple yet effective way to parameterize the density matrix composed of linear combination of irreducible tensor operators [43, 44, 45]
@@
where $T^{L}_{M}$ are the irreducible tensor operators complying with a trace rule such that Tr[${\rm Tr}\left[T_{M}^{L}(T^{L}_{M})^{\dagger}\right]$ = 3. It should be noted that a sum over the indices $L=1,2$ and $-L\leq M\leq L$. These tensor operators are usually formed from linear combination of the spin operators for spin-1 particles. One may refer to paper [24] for the exact form of these operators. The coefficients $A^{i}_{LM}$ and $C_{L_{1}M_{1}L_{2}M_{2}}$ are representing both the polarizations and spin correlations of pair of Z bosons. These coefficients are extracted from simulated or actual data, which is very challenging because both the CM frame of the colliding beam and the rest frame of the decaying particles should be established correctly so that the angular distributions of the final lepton can be determined precisely. Consequently, there will be multiple Lorentz boost or transformation between coordinate systems. In order to extract the coefficients of the density matrix, a relation between these coefficients and a direct observable in the experiment is needed. This has been done through a quantum state tomography for weak decays of massive bosons [46]. The high energy collider experiments directly provide the general observable quantity, cross section, which can be expressed in terms of decaying matrix density of the massive bosons. For example, the differential cross section for $ZZ\to\ell_{1}^{+}\ell_{1}^{-}\ell_{2}^{+}\ell_{2}^{-}$ can be written [45]
@@
where the solid angles $d\Omega^{\pm}\sin\theta^{\pm}d\theta^{\pm}d\phi^{\pm}$ are given in terms of the spherical coordinates for the momenta of the final charged leptons with respect to the rest frame of the decaying Z boson. $\rho_{V_{1}V_{2}}$ is the density matrix for the entangled $ZZ$ state explained previously. The decaying density matrix of a Z boson into charged leptons is given as [45, 24, 23]
@@
Now the coefficients in Eq. 9 can be obtained by integrating the above differential cross section over the solid angle of the leptons and using orthogonal property of the spherical harmonics:
@@
It is worth noting that the $ZZ$ system is in the singlet state because the third component of the spin along the boson momentum direction is conserved. This imposes strong constraints on the form of the density matrix, including only nine non-zero elements with the relation
@@
In the end, in this tensor-parametrization form of the density matrix, we are left with two free parameters that can be be determined from data.
@@ II.2 Quantum entanglement and Bell inequality for two-qutrit system
To construct the optimized Bell inequality, two observers A and B, each having two measurements—$A_{1}$ and $A_{2}$ for $A$, and $B_{1}$ and $B_{2}$ for $B$—are considered. For each value of $A$ and $B$, there are three possible outcomes. One can denote by $P(A_{i}=B_{j}+k)$ the probability that the outcomes $A_{i}$ and $B_{j}$ differ by $k$ modulo 3. Then we can obtain a simple function by linear sum of these functions:
@@
Once we have determined the coefficients given in Eq. 15 from the angular distributions of the leptons, one can immediately calculate the density matrix for the composite system, allowing to probe the entanglement and test Bell inequality violation through the expectation values of the Bell operator.
@@
where $\mathcal{B}$ is the quantum Bell operator [47] for spin-1 massive gauge boson. The construction of this Bell operator is always based on the linear combination of the three spin-1 component operators or the irreducible tensors [28, 26, 24]. It is worth noting that the violation of  17 obtained with the expectation values of the Bell operator may not be the maximal one because in the non-relativistic limit the joint $ZZ$ system is not totally in the singlet state (but close) [24]. Furthermore, we choose the outgoing direction of the on-shell boson as the polarization axis. However, a unitary rotation on the spin states may maximize the expectation value with following transformation on the Bell operator
where U and V are the independent three-dimensional arbitrary unitary matrices. Here we make use the Bell operator provided in [24] with the above unitary transformation: